# Estimating the Lensing Probability for Binary Black Hole Mergers in AGN disk by Using Mismatch Threshold

### v1

@@ III.1 Mismatch Arising from Lensing
Figure 3: Contour plot of the mismatch between lensed and unlensed waveform templates as a function of inclination angle and orbital phase. The solid black circle line represents mismatch values for $\rm{SNR}$ of $5,6,7,8$ and $9$, with $M_{\rm{Lz}}=10^{8}M_{\odot}$ and $a_{0}=500R_{\rm{S}}(M_{8})$ (left) or $a_{0}=5000R_{\rm{S}}(M_{8})$ (right). $Left$: The red solid circle line indicates the mismatch corresponding to the Einstein criterion, while the red dashed-dotted circle line shows the mismatch threshold for the minimum $\rm{SNR}$ of $4$. The red dotted circle line represents the threshold for the maximum $\rm{SNR}$ of $10$. The white solid horizontal line corresponds to the orbital phase at the Einstein criterion for $\iota=90^{\circ}$, $a_{0}=500R_{\rm{S}}(M_{8})$ and $M_{\rm{Lz}}=10^{8}M_{\odot}$. The white dashed-dotted horizontal line marks the minimum orbital phase for SNR = 4, and the white dotted horizontal line represents the maximum orbital phase for $\rm{SNR}=10$, both for $\iota=90^{\circ}$, $a_{0}=500R_{\rm{S}}(M_{8})$ and $M_{\rm{Lz}}=10^{8}M_{\odot}$. $Right$: The right panel is obtained by increasing the semi-major axis from $500R_{\rm{S}}(M_{8})$ (left panel) to $5000R_{\rm{S}}(M_{8})$, with all other parameters kept identical.
The calculated mismatch is shown in Figure 3, where we fix the semi-major axis to be $a=500R_{\rm S}$ (left panel) and $a=5000R_{\rm S}$ (right panel), and set the black hole mass to $10^{8}M_{\odot}$. We investigate how the mismatch varies with inclination angle $\iota$ and orbital phase $\phi$. It can be seen that the regions exhibiting significant differences between the lensed and unlensed waveforms are concentrated in a central circular area, corresponding to the well-aligned configurations. According to Equation (16), the SNR of the GW source determines whether the detector can distinguish between the lensed and unlensed waveforms, and thus affects the mismatch threshold. Based on this, we mark the contours corresponding to SNR values of 5, 6, 7, 8, and 9 with black circles in the plot. As the SNR increases from 5 to 9, the parameter space where the lensing effect can in principle be distinguished also expands. The dash-dotted red circular contour in the figure corresponds to $\epsilon_{\rm th}$ at an SNR of 4. This corresponds to a threshold orbital phase of $\phi_{\rm th}=4.55^{\circ}$. The dotted red circular contour represents $\epsilon_{\rm th}$ at an SNR of 10, corresponding to a threshold orbital phase of $\phi_{\rm th}=9.72^{\circ}$. The solid red circular marks the Einstein radius. It is seen that the lensing region determined by the mismatch threshold is generally larger than the Einstein radius. The SNR governs the distinguishable parameter space between lensed and unlensed waveforms, with higher SNR values corresponding to broader regions of resolvable differences. The intersection points between the vertical solid white line and the horizontal lines (solid, dashed, and dotted) indicate the maximum values of $\phi_{\rm th}$ when $\iota=90^{\circ}$ (the maximum threshold orbital phase is denoted as $\phi_{\rm th,max}$). Taking into account the symmetry of the point-mass lens, we will use these values to estimate the effective area in parameter space where lensing events may occur. The right panel of Figure 3 shows the case with a semi-major axis of $a=5000R_{\rm S}$, with all other parameters set identically. It is evident that increasing the semi-major axis significantly reduces the region where lensing may occur. The influence of the semi-major axis and orbital phase on the mismatch is shown in Figure (4)
Figure 4: The mismatch $\epsilon$ between lensed and unlensed waveform templates as a function of semi-major axis and threshold orbital phase $\phi_{\rm th}$. The red solid line indicates the mismatch corresponding to the Einstein criterion, while the red dashed-dotted line shows the mismatch threshold for the minimum $\rm{SNR}$ of $4$. The red dotted line represents the threshold for the maximum $\rm{SNR}$ of $10$. $\iota=90^{\circ}$ and $M_{\rm{Lz}}=10^{8}M_{\odot}$ are adopted here.
@@
Similar to Figure 3, we fix $\iota=90^{\circ}$ and set the mass of the SMBH to $10^{8}M_{\odot}$, and compute the variation of the mismatch threshold as a function of the semi-major axis and the orbital phase. The results are shown in Figure 4. It can be seen that, on the one hand, the region where the lensing effect is significant is concentrated around $\phi=0^{\circ}$. As the range of $\phi$ extends outward from $0^{\circ}$ towards both sides of the axis, the BBH gradually deviates from the line of sight, resulting in the decrease of mismatch. On the other hand, as the semi-major axis increases from $500\mathrm{R_{S}}$ to $5000\mathrm{R_{S}}$, the $\phi_{\rm th}$ (related to $\epsilon_{\rm th}$) range in which the lensing effect is prominent gradually shrinks. This is because when $\iota=90^{\circ}$, this corresponds to an edge-on disk. The angular position of the source is
@@
in this configuration, where $r=a/R_{\rm S}$. It is apparent that a larger semi-major axis imposes stricter geometric constraints on the system, reducing the available parameter space in exchange for the same lensing features (see Equation 18). As expected, the distinguishable lensing region in the right panel is smaller than that in the left panel. The contour lines in Figure 4 carry the same meaning as the corresponding line styles in Figure 3.
If we assume that the AGN disk is circular with an inclination angle of $90^{\circ}$, then, combined with the symmetry of the point-mass lens potential, rotating the AGN disk around the line of sight once would bring all GW sources with orbital phases smaller than $\phi_{\rm{th,max}}$ into configurations where lensing characteristics are distinguishable. The parameter $\phi_{\rm{th,max}}$ defines a spherical cap centered on the SMBH at the core of the AGN. We have illustrated the geometric configuration of this spherical cap in Figure 1 (b). Based on this geometry, we define the lensing probability as the ratio of the solid angle of the spherical cap to the total solid angle $4\pi$. The solid angle of the spherical cap is given by [66]
@@
$p({\rm lensing}|\phi_{\rm th,max})=\frac{1}{2}(1-\cos{\phi_{th,max}}).$ (20)
III.2 Lensing Probability
@@
In previous work, the Einstein radius is employed as a criterion (as the threshold) to evaluate the significance of lensing effects, serving as a boundary for calculating the probability of lensing events occurring [44, 67, 45]. In this simple estimation, the Einstein criterion requires that $\beta\leq\theta_{\rm Ein}$. From Equations (4) and (17), this leads to the inequality $r\cos^{2}\phi_{\rm Ein}+2\cos\phi_{\rm Ein}-r\geq 0$. Solving this inequality yields the threshold orbital phase corresponding to the Einstein criterion: $\phi_{\rm Ein}=\arccos{(\sqrt{r^{2}+1}-1)/r)}$. Substituting this result into Equation (20), and defining $x=(\sqrt{r^{2}+1}-1)/r)$, we obtain $p(\mathrm{lensing}|r)=(1-x)/2$, which resembles the result of Equation (A15) in [45]. For the purpose of comparison with our results. The orbital phase distribution related to the Einstein criterion as a function of semi-major axis is shown as a blue dashed line in Figure 5. A schematic illustration of the solid angle defined by the Einstein criterion is also shown in Figure 1 (b). The probability distribution related to the Einstein criterion as a function of semi-major axis is illustrated by the blue dashed line in Figure 6.
However, the Einstein radius serves only as an approximate boundary for the lensing effect. Further study indicates that the effective lensing region is generally larger than the Einstein radius [68, 69]. [60] indicates that the distinguishability of two GW waveforms is determined by the SNR, which in turn depends on the sensitivity of the detector. As seen from Equation (16), a higher SNR results in a smaller mismatch threshold. This implies that, even though there may be only slight differences between the template waveform and the lensed waveform, we may still have the opportunity to distinguish them through matched filtering for sources with a high SNR.
In the following, we employ Equation (20) to evaluate the lensing probability. Among the 93 sources observed by LVK to date [9, 10, 11, 12, 13] , only two of them have an SNR below 5, one source has an SNR greater than 30, and the majority of sources have an SNR exceeding 10. Based on this observational fact, we selecte two typical values of the SNR as 8 and 15, which related to the mismatch threshold of $\epsilon_{\rm th}\approx 0.008$ and $\epsilon_{\rm th}\approx 0.002$. For the case with a SNR of $\approx 8$ (and therefore $\epsilon_{\rm th}\approx 0.008$), we utilize aLIGOaLIGOO3LowT1800545 [11] as our power spectral density (PSD), and employ aLIGOAdVO4IntermediateT1800545 [61] as our PSD for SNR $\approx 15$ (and therefore $\epsilon_{\rm th}\approx 0.002$). For the lens parameters, we chose an edge-on disk ($\iota=90^{\circ}$). The semi-major axis of the source ranges from $500R_{\rm S}(M_{\rm Lz})$ to $2000R_{\rm S}(M_{\rm Lz})$. Based on these parameters, we calculated the mismatches for $\epsilon_{\rm th}(\rho=8)=0.008$ and $\epsilon_{\rm th}(\rho=15)=0.002$ and their related mismatch threshold orbital phase, respectively.
@@
Figure 5: The $\phi_{\rm th,max}$ distribution as a function of semi-major axis $a$ with $\epsilon_{\rm th}=0.002$ and $0.008$. All of the three lines are set $a=1000R_{\rm S}$ and $\iota=90^{\circ}$. The blue dashed line represents the Einstein criterion, which relates to the results of [45]. The blue solid line, black dashed line and gray dashed line are the power law fitting result of $\phi_{\rm th,max}\propto a^{-1/2}$.
Figure 6: The lensing probability distribution as a function of semi-major axis $a$ with $\epsilon_{\rm th}=0.002$ for $\rho=15$ and $0.008$ for $\rho=8$. All of the parameters are the same as in Figure 5. The blue dashed line represents the probability given by Einstein criterion, which relates to the results of [45]. The blue solid line, the black dashed line and the gray solid line are the power law fitting results of $p(lensing|\phi_{th,max})\propto a^{-1}$.
The relationship between the orbital phase $\phi_{\rm th,max}$ and semi-major axis $a$ is shown in Figure 5. The relationship between $\phi_{\rm th,max}$ and $a$ follows a power law of $\phi_{\rm th,max}\propto a^{-1/2}$. As expected, $\phi_{\rm th,max}$ gets smaller when $a$ increases. In Figure 1 $(b)$, we illustrate the cone angles corresponding to the maximum threshold orbital phase derived from the mismatch at $1000R_{\rm S}$, as well as the cone angles corresponding to the Einstein criterion. The cone angle represented by the black solid line and gray solid line corresponds to mismatch values of $\epsilon_{\rm th}=0.008$ and $\epsilon_{\rm th}=0.002$, respectively, while the cone angle in blue solid line is given by the Einstein criterion. Their corresponding probability distributions are given by Equation (20), as shown in Figure 6. The black solid line represents the case of $\epsilon_{\rm th}=0.008$, the gray solid line represents the case of $\epsilon_{\rm th}=0.002$, while the blue dashed line represents the Einstein criterion (equivalent to the mismatch threshold of $\epsilon_{\rm Ein}=0.06$, i.e. corresponding to the results of [45]).
Our results indicate that, for $a=1000R_{\rm S}$, the threshold orbital phase $\phi_{\rm th,max}$ and lensing probability calculated with the mismatch threshold $\epsilon_{\rm th}=0.008$ for $\rho=8$ (or $\epsilon_{\rm th}=0.002$ for $\rho=15$) will be $\sim 2$ (or $\sim 3$) and $\sim 6$ ($\sim 14$) times larger than the result with the Einstein criterion, respectively. The relationship between $p(lensing|\phi_{\rm th,max})$ and $a$ follows a power law of $p(lensing|\phi_{\rm th,max})\propto a^{-1}$. For the case of $a=500R_{\rm S}$, the orbital phase $\phi_{\rm th,max}$ calculated with the mismatch threshold $\epsilon_{\rm th}=0.008$ for $\rho=8$ (or $\epsilon_{\rm th}=0.002$ for $\rho=15$) will be $\sim 3$ (or $\sim 4$) times larger than the result with the Einstein criterion. We can see that, as the semi-major axis increases, the subthreshold orbital phase decreases. It is found that, if we take the semi-major axis of the source in units of Schwarzschild radius $R_{\rm S}$, the position of the source will not depend on the lens mass.
IV Conclusion and diskussion
@@
In this study, we propose that the detection criterion for gravitational lensing should be based on the PSD of the detector and the SNR of GW sources. We calculated the lensing probability for BBH mergers in AGN disks through systematic analysis of waveform-template mismatches between lensed and unlensed waveform template banks. For the current observational runs, the typical SNR values are $\sim 8-15$. In these cases, the orbital phase given by the mismatch threshold is $\sim 2-3$ times larger than previous estimates (i.e., predicted by the Einstein criterion). The lensing probability will be roughly 10 times larger than the previous result.
GW190521 was reported in [70]. Its plausible optical electromagnetic counterpart to a binary black hole merger was reported in [71]. As argued in [45], with the lensing probability given by the Einstein criterion, if assuming that all BBHs are GW190521-like, located in the migration trap at $a\simeq 331R_{S}$, one can find that with $\sim 100$ GW events, the probability of observing at least one lensed event is $14\%$. However, this percentage will be $\sim 60\%$ (or $\sim 88\%$) for the mismatch threshold $\epsilon_{\rm th}=0.008$ ($\rho=8$) (or $\epsilon_{\rm th}=0.002$ for $\rho=15$) if adopting our revised lensing probability. On the other hand, the non-detections will place more strict constraints on the fraction of AGN disk BBHs and even the birthplaces of BBH mergers. We might need to comment that our enhanced rate is ideal. In reality, the enhancement factor should be smaller when considering the real detections.
@@
As illustrated by the lensing configuration in Figure 1 ($a$), when $\iota=90^{\circ}$, a smaller $\phi$ corresponds to more pronounced lensing features, making it easier to identify the presence of lensing features. However, in this case, the lensing probability decreases (see Equation (20)). Conversely, as $\phi$ increases, the lensing features become less significant, but the lensing probability increases. This poses a significant challenge for the identification of lensed GW events. For the case of subthreshold searching [72], the amplitude of the second image is significantly de-magnified. This may lead to an inaccurate parameter estimation of the source [73, 74], and the weaker signal from the second image may be buried in the noise and therefore overlooked [46]. To justify our revised lensing probability, a dedicated subthreshold targeted search for lensing GW should be performed in the future. We stress that this work does not address the identification of lensed gravitational-wave signals. For studies on the identification of lensed GW events, we refer the reader to [55].
@@
GW lensing is also possible for BBHs in a dense star cluster. Therefore, distinguishing the physical origin of BBHs (in a star cluster or AGN disk) through GW lensing alone remains challenging. To further constrain the specific formation channel, one needs to consider the other detectable effects of the surrounding environment on the GW sources.

### v2

@@ III.2.1 Mismatch distribution: with O3 PSD
Figure 4: LeftLeft: Contour plot of the mismatch between lensed and unlensed waveform templates as a function of inclination angle and orbital phase. The solid, dashed, and dash-dotted circle contour line represents mismatch values for $\rm{SNR}$ of $5,6,$ and $10$, with $M_{\rm{Lz}}=10^{8}M_{\odot}$ and $a=500R_{\rm{S}}(M_{8})$. The dotted circle contour line in this figure indicates the mismatch corresponding to the Einstein criterion. RightRight: Mismatch contours as a function of orbital phase and semi-major axis for a fixed inclination $\iota=90^{\circ}$. The contour levels are the same as in the left panel.
In the left panel of FIG. 4, we fix the semi-major axis to $a=500R_{\rm S}$ and set the lens mass to $10^{8}M_{\odot}$. We investigate how the mismatch varies with inclination angle $\iota$ and orbital phase $\phi$. It can be seen that the regions exhibiting significant differences between the lensed and unlensed waveforms are concentrated in a central circular region, corresponding to the well-aligned configurations. This significant mismatch is caused by the superposition of two images. Given the extremely low probability of such a configuration (for $\phi\approx 0.05^{\circ}$, the probability is $\approx 10^{-7}$), we did not pursue further investigation in this work. For readers interested in the case of $\iota\approx 90^{\circ}$ and $\phi\approx 0^{\circ}$, we refer them to Ref. [44] for more details. According to Eq. (16), the SNR of the GW source determines whether the AGN lensing induced waveform deviation can be distinguished from the unlensed one, and thus affects the mismatch threshold. Based on the SNR distribution in FIG. 3, we mark the contours corresponding to SNR values of 5, 6, and 10 in FIG. 4. These SNR values correspond to the O3 PSD case, with $m_{1}=m_{2}=20M_{\odot}$ and source distances of $5000\rm{Mpc}$, $3000\rm{Mpc}$, and $1000\rm{Mpc}$, respectively.
The dash-dotted circular contour line in FIG. 4 corresponds to $\epsilon_{\rm th}$ at an SNR of $\sim 10$. This corresponds to a threshold orbital phase of $\phi_{\rm th}\approx 12^{\circ}$. The dashed circular contour line represents $\epsilon_{\rm th}$ at an SNR of $\sim 6$, corresponding to a threshold orbital phase of $\phi_{\rm th}\approx 9^{\circ}$. The solid circular contour line represents $\epsilon_{\rm th}$ at an SNR of $\sim 5$, corresponding to a threshold orbital phase of $\phi_{\rm th}\approx 7^{\circ}$. The dotted circular contour line marks the Einstein radius related mismatch value. It is seen that the AGN lensing region determined by the mismatch threshold is generally larger than the Einstein radius. The SNR governs the distinguishable parameter space between lensed and unlensed waveforms. As the SNR increases from 5 to 10, the parameter space where the lensing effect can be distinguished in principle also expands. We will subsequently use the phase angle $\phi_{\rm th}$ corresponding to the mismatch threshold to delineate the region where the AGN lensing effects are distinguishable. This region governs the AGN lensing probability.
@@
In the right panel of FIG. 4, we fix $\iota=90^{\circ}$ and set the mass of the SMBH to $10^{8}M_{\odot}$ and compute the variation of the AGN lensing induced mismatch as a function of the semi-major axis $a$ and the orbital phase $\phi$. The contour lines in the right panel of FIG. 4 carry the same meaning as the corresponding line styles in the left panel. It can be seen that, on the one hand, the region where the lensing effect is significant is concentrated around $\phi=0^{\circ}$. As the range of $\phi$ extends outwards from $0^{\circ}$ towards both sides of the axis, the BBH gradually deviates from the line of sight, resulting in a decrease in mismatch. On the other hand, as the semi-major axis increases from $500R_{\rm S}$ to $5000R_{\rm S}$, the $\phi_{\rm th}$ (related to $\epsilon_{\rm th}$) range in which the lensing effect is prominent gradually shrinks. This is because when $\iota=90^{\circ}$, this corresponds to an edge-on disk. The angular position of the source is
@@
in this configuration, where $r=a/R_{\rm S}$. It is apparent that a larger semi-major axis imposes stricter geometric constraints on the system, reducing the available parameter space in exchange for the same lensing features (see Eq. (19)).
If we assume that the AGN disk is circular with an inclination angle of $90^{\circ}$, then, combined with the symmetry of the point-mass lens potential, rotating the AGN disk around the line of sight once would bring all GW sources with orbital phases smaller than $\phi_{\rm{th,max}}$ into configurations where lensing characteristics are distinguishable. The parameter $\phi_{\rm{th,max}}$ defines a spherical cap centered on the SMBH at the core of the AGN. We have illustrated the geometric configuration of this spherical cap in FIG. 1 (b). Based on this geometry, we define the AGN lensing probability as the ratio of the solid angle of the spherical cap to the total solid angle $4\pi$. The solid angle of the spherical cap is given by Ref. [76]
@@
$p({\rm lensing}|\phi_{\rm th,max})=\frac{1}{2}(1-\cos{\phi_{\rm{th,max}}}).$ (21)
III.2.2 Mismatch distribution: with Aplus and ET PSD
In this subsection, in order to investigate the influence of different PSDs, we calculate the mismatch distributions for both the Aplus PSD and ET PSD scenarios, based on the results of Section III.1 and following the same procedures and parameter settings as used for the O3 PSD case. For the Aplus PSD, the source with the $20M_{\odot}+20M_{\odot}$ mass combination yields an SNR of approximately 20 at $1000\rm{Mpc}$, 10 at $3000\rm{Mpc}$, and 8 at $5000\rm{Mpc}$. According to Eq. (16), we calculate the threshold mismatches corresponding to these SNRs, which are represented by the contour lines in FIG. 5. The solid, dashed, and dash-dotted circle contour line represents mismatch values for $\rm{SNR}$ of 8, 10,8,\ 10, and $20$, with $M_{\rm{Lz}}=10^{8}M_{\odot}$ and $a=500R_{\rm{S}}(M_{8})$. For the ET PSD, the same mass combination yields an SNR of approximately 50 (25, 20) at $1000\rm{Mpc}$ ($3000\rm{Mpc}$, $5000\rm{Mpc}$). The corresponding threshold mismatches are indicated by the contour lines in FIG. 6.
These results demonstrate that at the same luminosity distance, higher detector sensitivity leads to a higher SNR. This higher SNRs are associated with smaller mismatch thresholds, which results in a larger distinguishable region for AGN lensing signatures, thereby increasing the AGN lensing probability. Consequently, we will adopt this mismatch threshold as the criterion for estimating the probability of AGN lensing in the following subsection.
Figure 5: The same as FIG. 4 but for Aplus PSD. LeftLeft: The solid, dashed, and dash-dotted circle contour line represents mismatch values for $\rm{SNR}$ of 8, 10,8,\ 10, and $20$, with $M_{\rm{Lz}}=10^{8}M_{\odot}$ and $a_{0}=500R_{\rm{S}}(M_{8})$. The dotted circle contour line in this figure indicates the mismatch corresponding to the Einstein criterion. RightRight: Mismatch contours as a function of orbital phase and semi-major axis for a fixed inclination $\iota=90^{\circ}$. The contour levels are the same as in the left panel.
Figure 6: The same as FIG. 4 but for ET PSD. LeftLeft: The solid, dashed, and dash-dotted circle contour line represents mismatch values for $\rm{SNR}$ of 20, 25,20,\ 25, and $50$, with $M_{\rm{Lz}}=10^{8}M_{\odot}$ and $a_{0}=500R_{\rm{S}}(M_{8})$. The dotted circle contour line in this figure indicates the mismatch corresponding to the Einstein criterion. RightRight: Mismatch contours as a function of orbital phase and semi-major axis for a fixed inclination $\iota=90^{\circ}$. The contour levels are the same as in the left panel.
III.3 Lensing Probability
In this subsection, we first calculate the variation of the threshold orbital phase and lensing probability with respect to the semi-major axis in AGNs, using the Einstein radius as the criterion. We then compute the variation of the threshold orbital phase and lensing probability with the semi-major axis under the O3, Aplus, and ET PSD scenarios, using the mismatch threshold as the criterion.
III.3.1 Einstein radius criterion
@@
In previous work, the Einstein radius is employed as a criterion (as the threshold) to evaluate the significance of lensing effects, serving as a boundary for calculating the probability of lensing events occurring [66]. In this simple estimation, the Einstein criterion requires that $\beta\leq\theta_{\rm Ein}$. From Eqs. (4) and (18), this leads to the inequality $r\cos^{2}\phi_{\rm Ein}+2\cos\phi_{\rm Ein}-r\geq 0$. Solving this inequality yields the threshold orbital phase corresponding to the Einstein criterion: $\phi_{\rm Ein}=\arccos{(\sqrt{r^{2}+1}-1)/r)}$. Substituting this result into Eq. (21), and defining $x=(\sqrt{r^{2}+1}-1)/r)$, we obtain $p(\mathrm{lensing}|r)=(1-x)/2$, which resembles the result in Ref. [66]. For the purpose of comparison with our results. The orbital phase distribution related to the Einstein criterion as a function of the semi-major axis is shown as a gray dotted line in the upper panel of FIG. 7. A schematic illustration of the solid angle defined by the Einstein criterion is also shown in FIG. 1 (b). The probability distribution related to the Einstein criterion as a function of the semi-major axis is illustrated by the gray dotted line in the lower panel of FIG. 7.
However, the Einstein radius serves only as an approximate boundary for the lensing effect. Further study indicates that the effective lensing region is generally larger than the Einstein radius [64]. Reference [71] indicates that the distinguishability of two GW waveforms is determined by the SNR, which in turn depends on the sensitivity of detectors and the intrinsic GW paremeters. We focus on the influence of detector sensitivity, the combination of source mass, and the source distance.
III.3.2 Mismatch threshold criterion
In the context of O3 PSD, with the mass combination of $20M_{\odot}+20M_{\odot}$ and other parameters of their default values, the SNR at source distances of $1000\rm{Mpc}$, $3000\rm{Mpc}$, and $5000\rm{Mpc}$ are 10, 6, and 5, respectively. Each of these SNRs determines a corresponding mismatch threshold. Subsequently, each mismatch threshold corresponds to a threshold orbital phase, which in turn determines the AGN lensing probability at the semi-major axis $a$ for an inclination angle of $90^{\circ}$.
@@ III.3.2 Mismatch threshold criterion
Figure 7: UpperUpper: The threshold orbital phase $\phi_{\rm th,max}$ distribution as a function of semi-major axis $a$ with source distanecs of $1000\rm{Mpc}$ (dash-dotted line), $3000\rm{Mpc}$ (dashed line), and $5000\rm{Mpc}$ (solid line). All of the three lines are set $\iota=90^{\circ}$. The gray dotted line represents the Einstein criterion, which relates to the results of Ref. [66]. The shaded line are the power law fitting result of $\phi_{\rm th,max}\sim a^{-1/2}$. LowerLower: Corresponding lensing probability as a function of semi-major axis. The shaded line are the power law fitting result of $p({\rm lensing|a})\sim a^{-1}$.
[…]
Figure 8: The same as FIG. 7, but for Aplus PSD.
Figure 9: The same as FIG. 7, but for ET PSD.
Following the same computational procedure as in the O3 PSD case, in FIGs. 8 and 9 we present the distributions of the threshold orbital phase and their AGN lensing probability as functions of the semi-major axis for the Aplus and ET PSD scenarios, respectively. Building on this, we have systematically investigated how different source distances and PSDs lead to variations in SNR, which subsequently determine the threshold orbital phase and ultimately influence the AGN lensing probability. Separately, we have also examined how different mass combinations and PSDs result in SNR differences, thereby affecting the threshold orbital phase and finally the AGN lensing probability. The corresponding results are presented in FIG. 10.
Figure 10: Probability as a function of semi-major axis. Three colors of purple, orange, and green represent O3 PSD, Aplus PSD, and ET PSD respectively, while three line types of solid, dashed, and dash-dotted correspond to three mass combinations: $m_{1}=m_{2}=20M_{\rm{\odot}}$; $m_{1}=35M_{\rm{\odot}},m_{2}=30M_{\rm{\odot}}$; $m_{1}=85M_{\rm{\odot}}$, $m_{2}=66M_{\rm{\odot}}$. The source distance are $D_{\rm S}=1000$ Mpc.
IV Conclusion and discussion
@@
On the aspect of AGN lensing probability, we propose that the criterion for gravitational lensing should be based on SNR. We calculated the lensing probability for BBH mergers in AGN disks through systematic analysis of waveform-template mismatches between lensed and unlensed waveform template banks. We summarize our results here.
We studied the effects of source distances. For O3 PSD, under fixed parameters, the AGN lensing probability based on the mismatch threshold is approximately a factor of $\sim 8$, $\sim 4$, and $\sim 3$ higher than that from the Einstein radius criterion at source distances of $1000\rm{Mpc}$, $3000\rm{Mpc}$, and $5000\rm{Mpc}$, respectively. Similarly, for Aplus PSD, the corresponding ratios are $\sim 15$, $\sim 8$, and $\sim 6$. Likewise, the corresponding ratios are $\sim 30$, $\sim 16$, and $\sim 13$ for ET PSD.
We studied the effects of BBH mass combination. For O3 PSD, the lensing probability based on the mismatch threshold is enhanced by factors of $\sim 10$ and $\sim 15$ for the $35+30M_{\odot}$ and $85+66M_{\odot}$ systems, respectively. Similarly, for Aplus PSD, the corresponding factors are $\sim 14$, $\sim 18$, and $\sim 26$ for the $20+20M_{\odot}$, $35+30M_{\odot}$, and $85+66M_{\odot}$ combinations. Likewise, the corresponding factors for ET PSD are $\sim 80$, $\sim 100$, and $\sim 150$ for the same mass combinations, respectively.
GW190521 was reported in Ref. [3]. Reference [66] indicated that, with the lensing probability given by the Einstein criterion, if assuming that all BBHs are GW190521-like and located in the migration trap at $a\simeq 331R_{\rm S}$, one can find that with $\sim 100$ GW events, the probability of observing at least one lensed event is $14\%$. In this work, we adopt the mismatch threshold as a criterion to calculate the AGN lensing probability for GW190521-like sources. The probabilities are found to be approximately $3\%$ for the O3 PSD, $6\%$ for the Aplus PSD, and $33\%$ for the ET PSD (Note that the source distance is set to $1000\rm{~Mpc}$; for a more realistic distance (e.g., $\sim 3300\rm{~Mpc}$), this percentage would be lower). Conversely, the non-detections will place more strict constraints on the fraction of AGN disk BBHs and even the birthplaces of BBH mergers. Meanwhile, this work presents the relationship between the lensing probability and the semi-major axis for BBHs in AGN disks. This relationship, when combined with the distributions of BBHs within the disk and the AGN population itself, may allow for a better estimation of the AGN lensing event rate.
We might need to comment that our enhanced rate is ideal. In reality, the percentage should be smaller when considering the real detections. First, waveform mismodeling [97] in both template generation and detector calibration can obscure subtle lensing signatures. Second, data quality issues may potentially affect lensing searches. For instance, GW200129 showed some significance in lensing searches [2], but in fact might have a significant glitch under subtraction [91]. Third, selection bias may also affect the inferred lensing probability. Reference [29] shows that, in current matched-filtering searches, lensed gravitational wave signals can lose SNR due to inaccurate parameter recovery and may fail signal-consistency tests, making them less detectable.
@@
As illustrated by the lensing configuration in FIG. 1 ($a$), when $\iota=90^{\circ}$, a smaller $\phi$ corresponds to more pronounced lensing features, making it easier to identify the presence of lensing features. However, in this case, the lensing probability decreases (see Eq. (21)). Conversely, as $\phi$ increases, the lensing features become less significant, but the lensing probability increases. This poses a significant challenge for the identification of lensed GW events. For the case of subthreshold searching [65, 35, 43, 39, 56, 54, 68, 51, 60, 77, 27, 83, 20, 67], where the signal falls below the detection threshold [2], the amplitude of the second image is significantly de-magnified. This may lead to an inaccurate parameter estimation of the source [82, 26], and the weaker signal from the second image may be sufficiently de-magnified as to fall below the detection threshold (i.e., become subthreshold) and consequently be missed by GW searching pipelines [11, 5, 6]. To justify our revised AGN lensing probability, a dedicated subthreshold targeted search for lensing GW should be performed in the future. We stress that this work does not address the identification of lensed gravitational-wave signals. For studies on the identification of lensed GW events, we refer the reader to Ref. [81].
@@
GW lensing is also possible for BBHs in a dense star cluster. Therefore, distinguishing the physical origin of BBHs (in a star cluster or AGN disk) through GW lensing alone remains challenging. To further constrain the specific formation channel, one needs to consider the other detectable effects of the surrounding environment on the GW sources[53], such as the Doppler effect [100], eccentricity effects [106, 105], and strong-gravity effects [119, 17, 88, 89].
Regarding the Doppler effect, Ref. [100] pointed out that the relative motion between the source and the lens can cause the observed GW signal from one image to generally appear blue- or redshifted compared to the GW signal from the other image. This velocity-induced differential Doppler shift gives rise to an observable GW phase shift between the GW signals from the different images. Regarding eccentric effects, the eccentricity of a binary’s orbit depends on the formation history of the binary. On the one hand, gravitational radiation causes the orbit of a BBH to circularize continuously [92, 113]; on the other hand, accretion and dynamical friction can potentially increase the eccentricity [117, 87, 105]. Therefore, when we are concerned with identifying the specific formation channel of a BBH, it may be more appropriate to focus on which radial locations exhibit more pronounced eccentric effects. Concerning strong-field effects, simulation results from Ref. [115] show that the location of migration traps can be as close as several tens of Schwarzschild radii. For BBHs trapped at these locations, strong-field effects are expected to become important [88]. Combining these effects with lensing to further investigate the formation channels of BBHs is a particularly interesting endeavor. We will explore these scenarios in future work.